OpenStudy (anonymous):

What is the simplified form of 15x^6/20y^5 times 6y^2/5x^4 ?

OpenStudy (anonymous):

Okay, so we just multiply the numerators together and the denominators together. \[{15x^6 \over 20y^5} {6y^2 \over 5x^4} = {90x^6y^2 \over 100x^4y^2}\]We reduce the fraction 90/100 to 9/10, and subtract the exponents for each variable. That is, the power of x is 6-4 = 2, and the power of y is 2-2 = 0. Then we have \[{9x^2y^0 \over 10} = {9x^2 \over 10}\]since anything to the 0 power is 1.